OpenStudy (anonymous):

cube root of (16x^5)? help and explain how to get the root of x^5?

OpenStudy (anonymous):

help me please?

OpenStudy (adamaero):

The regular square root is the squared root of... This can also be expressed a different way: \[\sqrt{x} = x ^{1/2}\] So you can guess what a cubed root means: \[\sqrt[3]{x} = x ^{1/3}\] or \[(bla-bla)^{1/3} = \sqrt[3]{bla-bla}\]

OpenStudy (adamaero):

As long as the stuff in the radical sign (or parenthesis)^to some number is multiplying, you can pull stuff out =)

OpenStudy (adamaero):

So \[?^{3} = 16\] and \[x^{5(1/3)} = ?\]

OpenStudy (anonymous):

@adamaero thanks!

OpenStudy (anonymous):

@adamaero would it be 2x (cube root)2x^2??